OpenStudy (anonymous):

Use the first three non-zero terms of the Maclaurin series to approximate the value of the definite integral

OpenStudy (anonymous):

\[\int\limits_{0}^{0.8} (1-cost)/t^2 dt\]

OpenStudy (anonymous):

you know the power series for cosine?

OpenStudy (anonymous):

And find an appropriate test to determine whether or not the series converges

OpenStudy (anonymous):

I forget :S

OpenStudy (anonymous):

man this is a pain. last one for me i think ok for cosine you only need to remember two things: \(\cos(0)=1\) so expansion starts with 1 and \(\cos(x)\) is even so you only have even terms. therefore \[\cos(x)=1-\frac{x^2}{2}+\frac{x^4}{4!}-\frac{x^6}{6!}+...\]

OpenStudy (anonymous):

subtract this mess from one and get \[\frac{x^2}{2}-\frac{x^4}{4!}+\frac{x^6}{6!}-...\]

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

now i guess we have to divide by \(x^2\) but that is easy enough. we get \[\frac{1}{2}-\frac{x^2}{24}+\frac{x^4}{720}\]three terms is apparently enough

OpenStudy (anonymous):

lord now we do what? take the anti derivative and plug in .8 and 0 and subtract. well at 0 we will get zero, but i have no idea what you get at .8

OpenStudy (anonymous):

again the anti derivative should be ok right?

OpenStudy (anonymous):

and the rest is an annoying calculator exercise whcih will leave up to you

OpenStudy (anonymous):

Thanks! Would you know what to use for this? find an appropriate test to determine whether or not the series converges

OpenStudy (anonymous):

series converges for sure, but i am not sure what test you are supposed to use. what do you have available?

OpenStudy (anonymous):

@Zarkon what does it mean "use appropriate test" for this one, any idea?

OpenStudy (anonymous):

Root test, ratio test, condensation test

OpenStudy (zarkon):

use a ratio test